Doaktown Products manufactures fishing equipment for recreational uses. The Miramichi plant produces the company’s two versions of a special reel used for river fishing. The two models are the M-008, a basic reel, and the M-123, a new and improved version. Cost accountants at company headquarters have prepared costs for the two reels for the most recent period. The plant manager is concerned. The cost report does not coincide with her intuition about the relative costs of the two models. She has asked you to review the cost accounting and help her prepare a response to headquarters.

Manufacturing overhead is currently assigned to products based on their direct labor costs. For the most recent month, manufacturing overhead was $328,800. During that time, the company produced 13,300 units of the M-008 and 2,900 units of the M-123. The direct costs of production were as follows.

M-008 M-123 Total
Direct materials $ 106,400 $ 116,000 $ 222,400
Direct labor 106,400 58,000 164,400

Management determined that overhead costs are caused by three cost drivers. These drivers and their costs for last year were as follows.

Activity Level
Cost Driver Costs M-008 M-123 Total
Number of machine-hours $ 139,300 8,000 2,000 10,000
Number of production runs 90,000 10 30 40
Number of inspections 99,500 25 25 50
Total overhead $ 328,800

Required:

a. How much overhead will be assigned to each product if these three cost drivers are used to allocate overhead? What is the total cost per unit produced for each product?

How much overhead will be assigned to each product if these three cost drivers are used to allocate overhead? What is the total cost per unit produced for each product? (Round "Total unit cost" to 2 decimal places.)

M-008 M-123
Total overhead      
Total unit cost

b. How much of the overhead will be assigned to each product if direct labor cost is used to allocate overhead? What is the total cost per unit produced for each product?

How much of the overhead will be assigned to each product if direct labor cost is used to allocate overhead? What is the total cost per unit produced for each product? (Round "Total unit cost" to 2 decimal places.)

M-008 M-123
Total overhead
Total unit cost      
Answer #1


Doaktown Products Requirement a М-008 М-123 Total Rate Direct materials 222400 106400 116000 Direct labor 106400 58000 164400
Answer #2

m-008

A) 8,000*13.93= 111,440

2,250*10= 22,500

25*1,990= 49,750

total cost (106,400+106,400)+183,690= 396,490

396,490/13,300= 29.8112781955 or 29.81


B) 425,600/13,300= 32